OpenStudy (anonymous):

At 8 am the Kankers left a campground, driving at 48 miles per hour. At 8:20 am, the Garcias left the same campground and followed the same route, driving at 60 miles per hour. At what time did they overtake the Smiths?

OpenStudy (anonymous):

d=vt We're looking for when they're at the same distance and we k now the Garcias left 20 minutes ahead of the Kankers. 48t=60(t-20) 48t=60t-1200 12t=1200 t=100 So the Kankers drove for 100 minutes and the Garcias drove for 80 minutes when they met. The Garcias overtook the Kankers at 9:40am